OpenStudy (anonymous):

Solve the differential equation dT/du =(T^2)(u^2)

OpenStudy (anonymous):

\[\frac{dT}{du}=T^2u^2\\ \frac{dT}{T^2}=u^2~du\] Integrate both sides, etc...

OpenStudy (anonymous):

I keep trying to do that but my answer does not match any of the ones that I have to choose from.

OpenStudy (anonymous):